A portion of Nichrome wire of radius 2.80 mm is to be used in winding a heating coil. If the coil must draw a current of 8.45 A when a voltage of 120 V is applied across its ends, find the following. (Take the resistivity of Nichrome to be 1.50 ? 10?6 ? · m.)

3. -10 points My Notes Ask Your Teacher A portion of Nichrome wire of radius 2.80 mm is to be used in winding a heating coil. If the coil must draw a current of 8.45 A when a voltage of 120 V is applied across its ends, find the following Take the resistivity of Nichrome to be 1.50 10-6 ?·m. (a) the required resistance of the coil (b) the length of wire you must use to wind the coil Need Help?Read It Submit AnswerSave Progress Practice Another Version
Answer #1

Given

Nichrome wire of radius r = 2.80 mm

I = 8.45 A

V = 120 V

R = ?

resistivity rho = 1.50*10^-6 ohm m

a) required resistance of the coil is  

V = I*R

R = V/I

R = 120/8.45 ohm

R = 14.20 ohm

c) we know that rho = R*A/l

l = R*A/rho

l = R*pi*r^2/rho

l = 14.20*pi*(2.80*10^-3)^2/(1.50*10^-6) m

l = 233.165 m

the length of the wire is l = 233.165 m